OpenStudy (anonymous):

Find the limit as t tends to 0 from the right. lim = (20*sqrt(t))/t t-0+

OpenStudy (badhi):

$$\frac{20\sqrt{t}}{t}=\frac{20}{\sqrt{t}}$$ when t approaches 0 from right side what do we get for $$\frac{1}{\sqrt{t}}$$